lottstodo85:

In rhombus ABCD, the measure of angle ABE equals 38°, find the measure of angle BAE.

lottstodo85:

76 104 90 52 are the answer choices

xxxFrancisco14xx:

Eliminate 76 and 90

xxxFrancisco14xx:

Take a guess 52 or 90

lottstodo85:

thxs

xxxFrancisco14xx:

which one do you think it is?

lottstodo85:

maybe 104

xxxFrancisco14xx:

52 is the answer, can I get best response

xxxFrancisco14xx:

oh wait I gave you a direct answer shoot

nevaehhh:

This was direct answer- but do you want me to explain it for you ?

xxxFrancisco14xx:

you can explain it I guess

xxxFrancisco14xx:

srry for the direct answer I forgot

nevaehhh:

Ok, so we see our angle measure 38... we need to add to each angle and see what we get. 76+38= ? 104+38= ? 90+38= ? 52+38=?

nevaehhh:

Once, we add all these, we get 114 142 128 90

lottstodo85:

oooooh that makes more sense

nevaehhh:

We were trying to find an angle measure that would add together to equal 90 degrees.

nevaehhh:

90 degrees=right angle

BraydenLove:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @nevaehhh This was direct answer- but do you want me to explain it for you ? \(\color{#0cbb34}{\text{End of Quote}}\) Lucky you're smexy